# On the Fujita Phenomenon for a Forced Spatio-Temporal Fractional Diffusion Equation

### v1

@@ 3. Proofs of the main results
To establish a contradiction, we proceed with a carefully chosen test function argument. We begin by introducing a smooth cutoff function $\psi\in\mathbf{C}^{2}(\mathbb{R}^{N})$ with the following properties
$0\leqslant\psi\leqslant 1\quad\text{and}\quad\psi(z)=\begin{cases}1&\text{if }|z|\leq 1,\\ 0&\text{if }|z|\geq 2.\end{cases}$
@@
By multiplying problem (1.1) by $\psi_{R}$, applying the integration by parts formula from Lemmas (2.5) and (2.4), we exploiting the properties of $\psi$, we derive the desired contradiction.
@@
$=\int\limits_{0}^{T}\int\limits_{\mathbb{R}^{N}}u(-\Delta)^{s}\psi_{R}(t,x)\,dx\,dt+\int\limits_{0}^{T}\int\limits_{\mathbb{R}^{N}}u\,\mathscr{D}_{t|T}^{\alpha}\psi_{R}(t,x)\,dx\,dt$
@@
Since $\psi$ is decreasing, the function $t\mapsto\psi\left(\frac{t}{R^{2\mathsf{s}/\alpha}}\right)$ is decreasing in $t$, hence $\mathscr{D}_{t|T}^{\alpha}\psi_{R}(t,x)\geq 0$ for $T$ sufficiently large. With $u_{0}(x)\geq 0$, the integrand is nonnegative, so
$\int\limits_{0}^{T}\int\limits_{\mathbb{R}^{N}}u_{0}(x)\,\mathscr{D}_{t|T}\psi_{R}(t,x)\,dx\,dt\geq 0,$
@@
We then have the estimate,

### v2

@@ 3. Proofs of the main results
To establish a contradiction, we proceed with a carefully chosen test function argument. We begin by introducing a smooth cutoff function $\psi\in\mathbf{C}_{0}^{\infty}([0,\infty))$ with the following properties
$0\leqslant\psi\leqslant 1\quad\text{and}\quad\psi(\tau)=\begin{cases}1&\text{if }0\leq\tau\leq 1,\\ 0&\text{if }\tau\geq 2.\end{cases}$
@@
By multiplying problem (1.1) by $\psi_{R}$, applying the integration by parts formula from Lemmas (2.1) and (2.5),we exploit the properties of $\psi$, we derive the desired contradiction.
@@
$=\int\limits_{0}^{T}\int\limits_{\mathbb{R}^{N}}u(-\Delta)^{s}\psi_{R}(t,x)\,dx\,dt+\int\limits_{0}^{T}\int\limits_{\mathbb{R}^{N}}u\,\mathscr{D}_{t|T}^{\alpha}\psi_{R}(t,x)\,dx\,dt.$
@@
Since $\psi$ is non-increasing, the function $t\mapsto\psi(t/R^{2s/\alpha})$ is non-increasing, and therefore $\mathscr{D}_{t|T}^{\alpha}\psi_{R}(t,x)\geq 0$ for $T$ sufficiently large. As $u_{0}(x)\geq 0$, the integrand is nonnegative, so
$\int\limits_{0}^{T}\int\limits_{\mathbb{R}^{N}}u_{0}(x)\,\mathscr{D}_{t|T}^{\alpha}\psi_{R}(t,x)\,dx\,dt\geq 0,$
@@
More precisely, we obtain
@@
4. Numerical simulation
In this section, we present numerical simulations illustrating the qualitative behavior predicted by Theorem 1.3. We focus on the one-dimensional fractional diffusion equation
$\partial_{t}^{1/2}u(t,x)+(-\Delta)^{1/4}u(t,x)=|u(t,x)|^{p}+e^{-|x|^{2}},\qquad(t,x)\in(0,\infty)\times\mathbb{R},$ (4.1)
which corresponds to the general model studied in this paper with $\alpha=\tfrac{1}{2}$, $s=\tfrac{1}{4}$ and $\sigma=0$. The forcing term $e^{-|x|^{2}}$ is nonnegative, localized in space, and satisfies $\int_{\mathbb{R}}e^{-|x|^{2}}\,dx>0$, so that all assumptions of Theorem 1.3 are fulfilled.
In this setting, the critical exponent defined in Theorem 1.3 reads
$p^{*}=\frac{N\alpha-2s\sigma}{N\alpha-2s(\alpha+\sigma)}=2,$
and therefore the theorem predicts the nonexistence of global weak solutions for $1<p<p^{*}$.
We numerically compare a subcritical case $p=1.5<p^{*}$ with a supercritical case $p=2.5>p^{*}$ in order to visualize the sharpness of this threshold.
Figure 1. Time evolution of $\|u(t,\cdot)\|_{L^{\infty}(\mathbb{R})}$ for $\alpha=\tfrac{1}{2}$ and $s=\tfrac{1}{4}$.
Figure 2. Spatial profiles $u(t,x)$ at representative late times for $\alpha=\tfrac{1}{2}$ and $s=\tfrac{1}{4}$.
Figure 1 displays the time evolution of the supremum norm $\|u(t,\cdot)\|_{L^{\infty}(\mathbb{R})}$. For $p=1.5<p^{*}$, the numerical solution grows rapidly and becomes unbounded in finite time, which is consistent with the blow-up (or nonexistence) mechanism established in Theorem 1.3. In contrast, for $p=2.5>p^{*}$, the solution remains bounded throughout the simulation, indicating a fundamentally different qualitative behavior.
Figure 2 complements this observation by highlighting the spatial mechanism associated with the blow-up. In the subcritical regime, the forcing induces a strong concentration near $x=0$, whose amplitude increases rapidly and leads to loss of boundedness. For $p>p^{*}$, the solution exhibits a much milder growth and no such concentration is observed within the simulated time range.
We stress that these numerical results do not constitute a proof of either blow-up or global existence. Rather, they provide a qualitative visualization of the critical threshold identified in Theorem 1.3. In particular, while the theorem guarantees nonexistence of global weak solutions for $1<p<p^{*}$ under the stated assumptions, it does not exclude blow-up for $p>p^{*}$ when the forcing is sufficiently large, nor does it assert global existence without additional conditions, such as the smallness assumptions appearing in Theorem 1.5.